### v2

# Dzyaloshinskii-Moriya interaction in Fe5GeTe2 epitaxial thin films

Abstract
Van der Waals ferromagnets, such as Fe5GeTe2, offer a promising platform for spintronic devices based on chiral magnetic textures, provided a significant Dzyaloshinskii-Moriya interaction (DMI) can be induced to stabilise the textures. Here, we directly measure DMI in epitaxial Fe5GeTe2 thin films using Brillouin light scattering spectroscopy and observe a consistent DMI ($D=0.04$ mJ/m2) across various thicknesses. Its weak thickness dependence, combined with the nominally symmetric film interfaces, suggests a bulk origin. Although we do not determine the microscopic mechanism, our findings are compatible with ab initio calculations linking DMI to partial ordering of Fe split sites. Additionally, we find a low magnetic dissipation ($\alpha<0.02$). The observed DMI, which could be further enhanced by optimising the Fe site ordering, combined with low dissipation, makes Fe5GeTe2 a strong candidate for exploring the dynamics of chiral magnetic textures in two-dimensional materials.
keywords:
van der Waals magnets, Dzyaloshinskii-Moriya Interaction, Chiral spin textures, Fe5GeTe2
Ferromagnetic van der Waals (vdW) materials present an interesting new class of materials for the study of nanomagnetism. Their structure, made of 2D layers with relatively weak inter-layer bonding, can produce very thin films with flat and clean surfaces, either by exfoliation from macroscopic crystals or epitaxial growth. This structure also lends itself naturally to the fabrication of heterostructures with good interfacial quality with a large variety of other vdW materials (semiconductors, metals, insulators, topological insulators, …) — ideal for studying interfacial magnetic effects that are the basis of spintronic devices. Among the growing list of vdW ferromagnets of interest to spintronics, the FexGeTe2 family stands out due to their high Curie temperature 1, metallic character, and large magneto-transport effects.
Moreover, there have been reports of chiral magnetic structures in FexGeTe2, opening the way to their use in the many proposed devices based on chiral textures such as skyrmions. Homo-chiral spin spirals were observed in exfoliated flakes of Fe3GeTe2 2 (thickness $t=185$ nm), and Néel skyrmions were reported in Fe3GeTe2 (70 nm)/[Co/Pd] 3 and WTe2/Fe3GeTe2 (32 nm) 4 heterostructures. In other members of the FexGeTe2 family, Néel skyrmions were observed in Fe5-xGeTe2 5, 6 (on 61 and 100 nm thick flakes), as well as in (Fe,Co)5GeTe2 (110 nm) 7 or in the related Fe3-xGaTe2 8, 9. On the other hand, other groups working also on flakes of Fe5GeTe2 observed only non-chiral Bloch bubbles 10, 11.
The observation of homo-chiral textures suggests a significant Dzyaloshinskii-Moriya interaction (DMI). Indeed, a large DMI parameter ($D=1.0$ mJ/m2) was estimated in ref. 4. A nonzero DMI requires a symmetry-breaking mechanism along the out-of-plane direction. However, both Fe3GeTe2 and Fe5GeTe2 are presumed to have inversion-symmetric crystal structures incompatible with DMI12. Nevertheless, several experimental studies suggest that the actual crystalline structure of FexGeTe2 may lack inversion symmetry. Possible sources of symmetry breaking include Fe split-site ordering 13, Fe self-intercalation in the van der Waals gaps 14, ordered Te vacancies 11, and asymmetric Fe vacancy distribution 15. Additionally, deviations from the expected vdW stacking sequence (ABCABC in Fe5GeTe2 and ABAB in Fe3GeTe2) could also break inversion symmetry.
DMI can also arise from asymmetric interfaces in thin magnetic layers. This mechanism has been well established in metallic systems, such as nanometric Co films in a AlOx/Co/Pt triplet film (up to $D=2$ mJ/m2 16, 17). Although this interfacial mechanism could be expected to be weaker in vdW materials due to their weak interlayer coupling, first-principle calculations by Park et al. 18 show that the interface of Fe3GeTe2 with its natural oxide can induce a significant DMI ($\approx 2$ mJ/m2 for a monolayer of Fe3GeTe2). Interfacial mechanisms, however, are only significant in ultra-thin systems, since their effects are progressively diluted with increasing thickness, while all the cited experimental observations involve relatively thick (30-200 nm) exfoliated flakes.
Characterizing the DMI solely from the observation of magnetic textures is challenging, as these are influenced by multiple interactions, including dipolar effects, magnetocrystalline anisotropy, and exchange. In some cases, dipolar interactions may even create seemingly chiral textures at the surface of thick samples 3, 19. To determine the origin of chiral magnetic textures and disentangle interfacial and bulk contributions, a more direct measurement of the DMI is essential. In this letter, we investigate the static and dynamic magnetic properties of epitaxial Fe5GeTe2 thin films of varying thickness using Brillouin light scattering (BLS) spectroscopy. By directly measuring the DMI, we unambiguously demonstrate the presence of a significant DMI of bulk origin. Furthermore, we show that the magnetic damping in these films is low for a metallic ferromagnet. These properties make Fe5GeTe2 a promising material for the study of the propagation of chiral magnetic textures.
Figure 1: a) Sample structure and diagrams of ordered split sites Fe5GeTe2 with broken centrosymmetry: full ordering (“DDD”) or 1/3 (“UUD”). Open circles represent empty sites. b) XRD spectrum for the three films. c) RHEED along the (100) and (110) reciprocal directions, obtained after growth (3-nm thick film). d) RHEED intensity profiles along the dashed line in c. The profiles are normalized to the (11) intensity and shifted vertically for clarity. Inset: Top view schematic of the “UUD” $\sqrt{3}\times\sqrt{3}$ reconstruction.
1 Results
1.1 Sample growth & structure
Fe5GeTe2 films ($t=$ 12, 8 and 3 nm) were epitaxially grown on Al2O3(0001) at 350°C by coevaporation of elemental Fe, Ge and Te, with a Fe:Ge flux ratio of 5 and with a slightly overstoichiometric Te flux, and then annealed $\sim 500$°C for 10 min. The 12-nm-thick film was characterised in depth in Ribeiro et al. 20, and all films were grown following the protocol reported in that work. The films were capped by 3 nm of Al deposited at room temperature, then naturally oxidized in air. This capping layer prevents the oxidation of Fe5GeTe2 and provides similar top and bottom interfaces (Fig. 1a).
The X-ray diffraction (XRD) specular scans (Fig.1b) show single-phase Fe5GeTe2 with a constant $c$-axis lattice parameter (0.98 nm per monolayer) across all thicknesses (see S.I.), consistent with the good crystalline quality previously observed for the 12-nm film by XRD and scanning transmission electron microscopy (STEM) 20. The RHEED patterns of Fe5GeTe2 (8 nm) in the (100) and (110) reciprocal directions, measured after growth, are shown in Fig. 1c. Similar anisotropic diffraction patterns are observed at all thicknesses and demonstrate the good epitaxy on sapphire. However, faint (10) diffraction rods are visible in the (110) direction (peak at $q/q(11)=1/\sqrt{3}$ in Fig. 1d). They denote the presence of a small proportion of domains that are rotated by 30° from the main crystal orientation. This feature is more apparent for the thinnest layer and disappears for the 12 nm film, indicating an improvement of the crystal in-plane orientation quality with increasing thickness.
In addition, the RHEED patterns display a ($3\times 1$) reconstruction characteristic of a ($\sqrt{3}\times\sqrt{3}$) R30° superstructure (peaks at q/q(11) = 1/3 and 2/3 in Fig. 1d). This superstructure is observed in all three films, and is more apparent at larger thicknesses. The nature of this feature will be discussed later.
Figure 2: a) $M_{s}(T)$ of the three films under in-plane field (SQUID measurement). b) LMOKE hysteresis loops of the three films close to 150 K. c) In-plane magnetic domains in the 12-nm-thick sample imaged by LMOKE (the contrast orientation is left/right).
1.2 Magnetic properties
The spontaneous magnetization ($M_{s}$) was measured by SQUID magnetometry versus temperature (Fig. 2). Thinner films showed lower $M_{s}$, with the 3-nm film showing about one-third of the $M_{s}$ of the 12-nm film at 150 K (see Table 1). The Curie temperature ($T_{C}$) of the thicker films (297–299 K) is similar to bulk values (270–310 K 1), whereas the 3-nm film had a lower $T_{C}$ (258 K). The evolution of $M_{s}$ and $T_{c}$ indicates that a thickness of a few layers is needed to fully develop the bulk magnetic properties of the films. The $M_{s}(T)$ curves showed a monotonic trend, unlike previous reports of various transitions in $M_{s}(T)$ 21, 10, 13, 22. However, those studies focused on virgin films, whereas our samples had been previously magnetized. These transitions may also be specific to bulk samples rather than epitaxially-grown films, suggesting structural or chemical differences between these two classes of materials.
[…]
$t$ $T_{C}$ $\mu_{0}M_{s}$ $B_{K}$ $g$ $\alpha$ $v_{DMI}$ $D$
nm K T T m/s µJ/m2
3 258 0.19 0.59-0.59 2.3 <0.08<0.08 $18\pm 4$ $22\pm 5$
8 299 0.36 0.84-0.84 2.2 <0.03<0.03 $20.5\pm 1.5$ $47\pm 3$
12 297 0.57 1.0-1.0 2.1 <0.02<0.02 $10.9\pm 1.3$ $41\pm 5$
Table 1: Material parameters at 150 K for the three films. The $T_{C}$ and $\mu_{0}M_{s}$ were determined from SQUID magnetometry, $B_{K}$ and the gyromagnetic g-factor were taken from fits of the BLS frequency versus field curves, the $\alpha$ (an upper bound; see text) from the peak width, and $v_{\rm DMI}$ from fits of the Stokes/anti-Stokes frequency difference. The negative $B_{K}$ represents an easy-plane anisotropy.
The samples were further characterised by studying the spin wave frequency as a function of field and wave vector using BLS spectroscopy (analysed using the laws of dispersion determined in refs. 23, 16, 24; see S.I.). The negative anisotropy field $B_{K}$ extracted from the BLS measurements of frequency versus field (Fig. S2) confirms the easy-plane (i.e., ab plane) anisotropy (at 150 K; Tab. 1). Its magnitude is significantly larger than the expected contribution from shape anisotropy (given by $\mu_{0}M_{s}$), a behavior previously observed in Fe5GeTe2 20, 21. This suggests either the presence of an intrinsic easy-plane magnetocrystalline anisotropy or that the magnetisation is inhomogeneous across the vdW layers and gaps, leading to a stronger shape anisotropy than expected from the measured $M_{s}$ (see S.I.).
The g-factor obtained from the BLS measurements consistently falls in the range g=2.12.3g=2.1-2.3, showing no significant temperature dependence, in close agreement with previous reports 25, 21. The notable deviation from the free-electron value of $g=2.0$ suggests a substantial orbital moment and significant spin-orbit coupling in the material. This is consistent with the orbital/spin moment ratio measured in ref. 20.
The Gilbert dissipation parameter ($\alpha$) can be inferred from the BLS peak width ($\Gamma$), similar to its extraction from ferromagnetic resonance measurements. However, $\Gamma$ is also broadened by inhomogeneities in the film’s magnetic properties, which are typically disentangled from the effect of $\alpha$ with high magnetic fields taking advantage of their distinct field dependence. Since the available field was insufficient for this separation, we estimated an upper bound, $\alpha_{\rm eff}$ (see Fig. S3 and the S.I.). The lowest $\alpha_{\rm eff}$ values for each film are listed in Table 1. The $\alpha_{\rm eff}$ is larger for thinner films, which could be due to a higher dissipation, as is often observed in other magnetic films (e.g. ref. 26). It may also be due to a higher inhomogeneity of the magnetic properties, as the properties of thinner films are more affected by variations of the thickness or of the interfaces. Overall, $\alpha_{\rm eff}$ as low as 0.02 were observed, lower than reported measurements in bulk Fe5GeTe2 21.
This low $\alpha$ is noteworthy, as metallic systems containing heavy elements typically exhibit large damping. This behavior is commonly interpreted using the breathing Fermi surface model 27, which predicts that $\alpha$ scales quadratically with the spin–orbit coupling (SOC) strength, linearly with the density of states (DoS) at the Fermi level, and with the electron relaxation time $\tau$ (itself proportional to the conductivity). The DoS of Fe5GeTe2 is reported to be comparable to that of iron 28. Conversely, its low conductivity suggests a short $\tau$. In addition, its weak magnetocrystalline anisotropy points to a modest SOC. Indeed, Te primarily bonds to the weakly magnetic Fe1 sublattice 20, limiting its contribution to damping. These factors may explain the observed $\alpha$, but they also highlight the need for an ab initio study of damping in these materials. Moreover, such low magnetic dissipation facilitates the motion of spin textures, such as skyrmions, further adding to the interest of Fe5GeTe2 as a material for Spintronics.
Figure 3: a) A BLS spectrum of the 12 nm film ($\phi=45$º, 150 K, $B=130$ mT). The green lines show the Lorentzian fit (fitted centres are written above the peaks). The central peak corresponds to non-scattered light and shows the instrumental width. b) Schematic of the BLS geometry. c) Stokes/anti-Stokes frequency difference, $\Delta f$, versus wave vector $k$ for the three films (points; taken at $B=$ 211, 250, and 130 mT, respectively.) The linear fits (lines) are used to extract $v_{\rm DMI}$ and $D$. d) Extracted $D$ versus sample thickness $t$ (points) and fits assuming a uniform bulk origin (constant blue line) and an interfacial origin ($\propto 1/t$, green line).
1.3 DMI
The Stokes and anti-Stokes peaks in the BLS spectrum (Fig. 3a) correspond to cycloidal spinwaves of opposite rotation sense (Fig. 3b), and are subject to an opposite energy contributions from the DMI. In the limit of thin films ($kt\ll 1$, where $k$ is the wave vector of the spinwave determined by the incidence angle of the light), as is the case here, the difference in frequency between the Stokes and anti-Stokes peaks ($\Delta f\equiv|f_{\rm anti-Stokes}|-|f_{\rm Stokes}|$) can be fully attributed to this effect, and can be used to determine the DMI parameter $D$ 29, 16:
$\Delta f=\frac{2\gamma}{\pi M_{s}}Dk,$ (1)
where $\gamma=g\mu_{B}/\hbar$ is the gyromagnetic ratio (and $\mu_{B}$ and $\hbar$ are the Bohr’s magneton and the reduced Planck constant, respectively.) As there may be a small constant $\Delta f$ due to an imperfect alignment of the spectrometer, several spectra were measured and the slope of $\Delta f(k)$ (which we dub $v_{\rm DMI}$) was determined. The expected reversal of the sign of $\Delta f$ with a reversal of the field direction was also verified. Fig. 3c shows the measured $\Delta f(k)$ for the 3, 8 and 12 nm films at 150 K. We observe that there is clearly a linear slope and thus a finite $D$ in all three films. All three show a slope, and thus a DMI, of the same sign. With the geometry of our setup ($\bm{\mathbf{k}}_{\rm Stokes}\times\bm{\mathbf{B}}$ pointing into the film surface, Fig. 3b), this positive slope corresponds to favoured right-handed spinwaves, and thus a positive DMI parameter. The calculated $D$ parameters (eq. 1) are shown in Tab. 1 for the three films. As a check, the DMI of the 8-nm film was also measured at 100 K and 300 K (Fig. S4 in S.I.), yielding a similar value at 100 K and no detectable DMI at 300 K, as expected near $T_{C}$.
2 Discussion
The measured DMI (0.02 to 0.05 mJ/m2) is comparable to values in many spintronic systems, though lower than the strongest known effects (e.g., $\sim$0.2 mJ/m2 for Co/Pt at 8 nm thickness 16). Micromagnetic simulations30 (see S.I.) show that this DMI suffices to induce chiral magnetic textures if the anisotropy is tuned to favour out-of-plane magnetisation.
The observation of DMI that favours cycloidal textures requires broken inversion symmetry along the $c$-axis, either in the bulk of the material or due to asymmetric interfaces. This symmetry breaking must persist over a large lateral scale, as BLS measurements average over the 30-µm laser spot and are consistent between points several 10s of µm apart. The dependence of $D$ on film thickness $t$ can help distinguish between bulk and interfacial origins: a bulk contribution would result in a constant $D$, whereas an interfacial contribution should scale as $D\propto 1/t$ 16. Figure 3d shows $D(t)$ alongside best-fit curves for both scenarios. Clearly, $D(t)$ deviates from the $1/t$ trend expected for interfacial DMI, in alignment with the rather electronically similar interfaces of the films (Al2O3/Fe5GeTe2 and Fe5GeTe2/AlOx). Instead, the $D(t)$ data is more compatible with the constant behaviour of a bulk origin, with a $D=0.04$ to 0.05 mJ/m2, as measured in the thickest films. The weaker DMI in the 3-nm-thick film, which correlates with lower $M_{s}$ and $T_{c}$, can be attributed to weaker ferromagnetism in the ultrathin regime.
[…]
Chakraborty et al. 15 attributed a possible bulk DMI in Fe3GeTe2 to intercalated Fe in the vdW gaps and Fe vacancies. In our films, however, intercalated Fe was not detected 20 (see Fig. S1 in S.I.); it would also raise the Curie temperature 14, which was not observed. Furthermore, Rutherford backscattering (RBS) analysis show a stoichiometric Fe content (Fe:Ge = 5.00±0.08) 20, indicating at most a very low concentration of Fe vacancies.
A second class of mechanisms involves symmetry breaking within the Fe5GeTe2 vdW layers. This is supported by our RHEED observations of a $\sqrt{3}\times\sqrt{3}$ superstructure, which cannot derive from a specific vdW stacking. A non-centrosymmetric phase with a $\sqrt{3}\times\sqrt{3}$ superstructure was reported by Gopi et al. 11, who ascribed it to a pattern of 1/3 Te vacancies ordered along the perpendicular direction. Since this phase represented only a small fraction ($10^{-2}$) of the studied bulk crystals, it was deemed to have a negligible effect on DMI. While this phase might be more prevalent in epitaxial films, RBS showed no significant Te depletion in our films 20 (Te:Ge = $2.0\pm 0.08$ for the 12-nm-thick film), suggesting that this mechanism is unlikely to explain our results.
Fe5GeTe2 is expected to have two partially-filled Fe sites (“split sites”) at different vertical positions within each monolayer, “D” and “U” (Fig. 1a). Several groups interpreted the $\sqrt{3}\times\sqrt{3}$ superstructure as the ordering of Fe split sites 13, 1. In the centrosymmetric structure, these sites are equally and randomly populated. However, ab initio calculations showed that the “UDD” and “UUD” configurations are energetically more favorable 33. While our RHEED measurements cannot determine whether this ordering extends beyond the top layer, electron microscopy analyses in Ref. 1 showed a similar short-range order in the bulk of Fe5GeTe2. Three groups have calculated the DMI assuming a perfectly ordered “UUU” arrangement in a monolayer, using different computational methods 34, 35, 5. Their predicted $|D|$ values vary significantly (0.76, 0.6, and 0.06 mJ/m2), despite all employing commonly used techniques. Our measured $|D|$ (0.04 – 0.05 mJ/m2) is comparable to the lowest theoretical estimate, and much lower than the highest prediction. This quantitative difference may be due to several reasons. First, the calculations do not account for temperature, which reduces the macroscopic $D$. Second, our films are not ordered in the “UUU” configuration since this would be inconsistent with the observed $\sqrt{3}\times\sqrt{3}$ superstructure. According to Ghosh et al. 35, the DMI of the “UUD” or “UDD” configurations is only slightly lower than that of the “UUU” configuration ($|D|=0.05$ instead of 0.06 mJ/m2), contrary to the naïve expectation of a 2/3 reduction. Finally, by symmetry, the “UUD” and “UDD” configurations contribute equally but with opposite sign DMI. While one of the two must be dominant to account for the observed DMI, the other may be present in a smaller amount and partially compensate the net DMI.
Assuming that the DMI arises from an ordering of the Fe split sites, the sign of $D$ can be used to infer the dominant split-site configuration. Both refs. 34 and 5 found that a fully-ordered “UUU” configuration favours counterclockwise spinwaves, corresponding to a negative DMI. Therefore, the observed positive DMI suggests that the predominant ordering in our films is “UDD.”
3 Conclusion
In conclusion, we have used BLS to quantitatively measure a significant DMI favouring cycloidal chiral textures in several epitaxial Fe5GeTe2 films with nominally symmetric interfaces. The DMI values were consistent across films of different thicknesses and exhibited the same sign, indicating a bulk origin for the DMI. Although our measurements cannot reveal the microscopic origin of the DMI, they are consistent with a partial ordering of the Fe split sites during film growth, a mechanism supported by published ab initio calculations showing that such ordering induces DMI.
These results confirm that Fe5GeTe2 can host a significant DMI and suggest that the reported magnetic textures in Fe5GeTe2 are likely chiral in nature. If the mechanism for the non-centrosymmetric structure is indeed induced during the growth as we suggest, the DMI may be enhanced by adjusting the growth conditions or by employing chemical substitutions which preferentially affect the split-site sublattice, such as with Co 36 or Ni 37. Additionally, DMI might be stronger in Fe5GeTe2 films synthesized or processed using other techniques. The bulk DMI mechanism observed here may coexist with interfacial contributions, such as those involving oxide layers as considered in previous studies, offering opportunities to further enhance the DMI.
With its significant bulk DMI, low magnetic dissipation, and the potential to engineer its interfaces, Fe5GeTe2 emerges as a promising material for exploring the dynamics of chiral magnetic textures and enabling spintronic applications.
{suppinfo}
The supporting information (S.I.) includes more details and figures relating to the experimental methods, the structural analysis of the films, the fitting and analysis of the BLS spectra, additional BLS measurements, and micromagnetic simulations of the effects of the measured DMI.
{acknowledgement}
The authors thank V. Jeudy for his support with cryogenic measurements and H. Okuno for the STEM image of Fig. S1. Magnetometry measurements were kindly performed by E. Rivière at ICMMO (CNRS) and at the Physical Measurements Platform at the LPS. This work was supported by the French National Research Agency (ANR) through the projects ELMAX (ANR-20-CE24-0015), ESR/EQUIPEX+ 2D-MAG (ANR-21-ESRE0025), NEXT (ANR-23-CE09-0034), the FLAG-ERA grant MNEMOSYN (ANR-21-GRF1-0005-01), the LANEF framework for mutualised infrastructure (ANR-10-LABX-51-01), and the France 2030 government investment plan (PEPR SPIN: SPINMAT ANR-22-EXSP-0007 and CHIREX ANR-22-EXSP-0002).